# Attention-Based Reading, Highlighting, and Forecasting of the Limit Order Book

### v1

@@ Abstract
Managing high-frequency data in a limit order book (LOB) is a complex task that often exceeds the capabilities of conventional time-series forecasting models. Accurately predicting the entire multi-level LOB, beyond just the mid-price, is essential for understanding high-frequency market dynamics. However, this task is challenging due to the complex interdependencies among compound attributes within each dimension, such as order types, features, and levels. In this study, we explore advanced multidimensional sequence-to-sequence models to forecast the entire multi-level LOB, including order prices and volumes. Our main contribution is the development of a compound multivariate embedding method designed to capture the complex relationships between spatiotemporal features. Empirical results show that our method outperforms other multivariate forecasting methods, achieving the lowest forecasting error while preserving the ordinal structure of the LOB.
@@
limit order book , multivariate time-series forecasting , attention mechanism , spatiotemporal embedding
[…]
[inst1]organization=Department of Statistics, Purdue University,city=West Lafayette, postcode=47907, state=IN, country=USA
@@ 1 Introduction
Given the complexities involved in predicting the full dynamics of a limit order book (LOB), many researchers have narrowed their focus to forecasting the mid-price, which is the average of the best bid and ask prices. Various studies have utilized different models for this task, including regression-based approaches (Ntakaris et al., 2018), support vector machines (Kercheval and Zhang, 2015), and deep neural networks (Sirignano, 2019; Zhang et al., 2019) to predict future mid-price movements. While mid-price serves as a useful indicator of overall market trends, Figure 1 highlights that the LOB can exhibit different features such as order depths at different price levels and bid-ask spreads (BAS), even when the mid-price remains the same. Therefore, relying solely on mid-price analysis may be insufficient when more detailed depth information is needed, such as in optimal execution problems, where a comprehensive understanding of the full LOB structure is essential.
@@
Attention-based models like Transformer (Vaswani et al., 2017) and BERT (Devlin et al., 2019) excel at sequence-to-sequence prediction by assigning varying weights to different parts of the data, enabling them to focus on the most relevant information for more accurate predictions. These models have proven highly effective in tasks such as natural language processing (Yang et al., 2021; Galassi et al., 2020; Cui et al., 2019), speech recognition (Zeyer et al., 2018; Chorowski et al., 2015; Bahdanau et al., 2016), and time series forecasting (Shih et al., 2019; Fan et al., 2019; Du et al., 2020). However, their application to limit order book forecasting remains limited, with notable examples including Arroyo et al. (2024) in survival analysis, Liu et al. (2024) using vision transformers, and Wu et al. (2024) employing wavelet decomposition.
In this study, we propose employing attention-based models, particularly those introduced in Spacetimeformer (Grigsby et al., 2022), to address the unique challenges posed by the spatiotemporal forecasting of limit order book. Our approach builds upon the ideas presented in this work and applies them specifically to the task of predicting limit order book movements. Our embedding approach integrates spatiotemporal features while preserving the structure of the LOB, leading to superior predictive accuracy compared to traditional forecasting methods. These advancements demonstrate the potential of attention-based models to enhance predictive performance in high-frequency trading environments.
2 Data and method
@@ 2.1 Data structure and preparation
We utilize the LOBSTER dataset (Huang and Polak, 2011), which offers real-world LOB data, including both message and order book information for selected stocks. We specifically focus on five major technology companies: AAPL, GOOG, INTC, MSFT, and AMZN. The sample covers the trading day of June 21, 2012, from the market opening at 9:30 am to the closing at 4:00 pm, capturing between 300,000 to 600,000 snapshots per stock, all recorded with millisecond time resolution at Level-5 of the order book. The order book file tracks the evolution of the LOB, while the message file logs the events that trigger updates, with timestamps ranging from milliseconds to nanoseconds.
@@
2.2 Stationary transformation and scaling
@@ 2.3 Compound multivariate embedding
$\text{Attention}(Q,K,V)\in\mathbbm{R}^{L_{x}\times d}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V,$ (3)
@@ 3 Results
Table 1: Prediction results for sample test data: Mean Squared Error (MSE) and Mean Absolute Error (MAE) are computed for each method across selected dimensions. The mid-price is determined by averaging the best bid and ask prices. Loss metrics are provided for price, volume, and overall performance on level-5 limit order book (LOB) data. Structure loss quantifies the violations of the ordinal structure, while total loss represents the weighted sum of the forecasting loss. All values are calculated after scaling.
@@
[…]
@@
    Entire refers to the sum of losses from all five levels of prices and volumes.
@@
We present the prediction results for sample test data, with a focus on various performance metrics discussed in Section 2.4. Our experiments compare various sequence-to-sequence prediction models of prior work, including a linear autoregressive model (Linear), a standard LSTM network (Schmidhuber et al., 1997) (LSTM), a transformer with temporal attention (Temporal) as introduced in Informer (Zhou et al., 2021), and a spacetimeformer model (Spacetime) that incorporates both spatial and temporal attention (Grigsby et al., 2022). We also evaluate the performance of our attention-based model with compound multivariate embedding (Compound) in comparison to the other techniques.
As shown in Table 1, our compound multivariate embedding method consistently ranked first or second among both time series and attention-based models, achieving the lowest forecasting error across the entire dataset. Notably, the compound multivariate embedding model significantly outperformed all others in structure loss—by a ratio of 0.2 to 0.6 times smaller—indicating far fewer violations of the ordinal structure in its predictions. This highlights the model’s superior ability to capture structural dependencies through the compound multivariate embedding process.
@@
Figure 6 and Figure 7 provide visual comparisons of the actual versus predicted price and volume movements for the level-5 LOB data. We selected the first window of the test period for comparison. In Figure 6, the price forecasting results demonstrate that the compound multivariate embedding method (B) closely tracks the downward trend of actual price movements (A), offering more accurate predictions than the temporal (C) and spatiotemporal (D) embedding methods. While both our approach and the spatiotemporal embedding effectively capture the overall ordinal structure of multi-level prices, the temporal embedding fails to preserve the ordinal relationships between dimensions. This aligns with the concepts of the embedding methods, as the temporal embedding processes multiple time series without incorporating variable embeddings, resulting in structural violations. Similarly, in Figure 7, the volume forecasting graphs show that the compound multivariate embedding method (B) yields overall volume distributions that more closely match the actual order sizes (A), outperforming the other attention-based models (C and D).
@@ 4 Conclusions
In this study, we addressed the challenge of forecasting multi-level limit order book (LOB) data, a task that often exceeds the capabilities of conventional time-series forecasting models. Our primary contribution is the development and implementation of a compound multivariate embedding method within advanced sequence-to-sequence models. This method effectively captures the complex interdependencies among various LOB attributes, including order types, features, and levels. The empirical results demonstrate that our approach consistently outperforms existing multivariate forecasting methods, achieving the lowest forecasting errors while maintaining the ordinal structure of the LOB. These findings highlight the effectiveness of our method for financial time-series forecasting, particularly in the context of high-frequency trading. As a direction for future work, our method could be extended to address optimal execution problems, with further optimization for real-time forecasting scenarios to enhance its practical applicability in dynamic trading environments.
Conflict of interest
The authors have no conflict of interest to disclose.
Data availability
The authors do not have permission to share data.

### v2

@@ Abstract
Managing high-frequency data in a limit order book (LOB) is complex, often beyond the capabilities of conventional time-series models. Accurate prediction of the multi-level LOB, not just the mid-price, is crucial for understanding market dynamics but is difficult due to the interdependencies among attributes like order types, features, and levels. This study introduces advanced sequence-to-sequence models to forecast the entire multi-level LOB, including prices and volumes. Our key contribution is a compound multivariate embedding method that captures spatiotemporal relationships. Empirical results show that our method outperforms others, achieving the lowest forecasting error while maintaining LOB structure.
@@
limit order book; multivariate time-series forecasting; attention mechanism; spatiotemporal embedding
@@ 1 Introduction
Another challenge in time series forecasting, including LOB prediction, is the non-stationarity of the data. Non-stationarity refers to changes in the statistical properties of a time series over time, such as mean, variance, or autocorrelation, which can significantly impact forecasting accuracy. Shifts in trends and structural changes over time can result in distributional differences between the training and test periods, undermining model performance. Techniques such as adaptive attention mechanisms Liu et al. (2022) and trainable normalization methods Kim et al. (2021) have been proposed to address these issues, but still struggle with abrupt changes in high-frequency data and require careful tuning to handle long-term dependencies.
Given the complexities of predicting the full dynamics of an LOB, many researchers have focused on forecasting the mid-price, defined as the average of the best bid and ask prices. Various machine learning models have been applied to this task, including regression-based approaches (Ntakaris et al., 2018), support vector machines (Kercheval & Zhang, 2015), and deep neural networks (Liu et al., 2024; Tsantekidis et al., 2020; Zhang et al., 2019) to predict future mid-price movements. The mid-price is a general market trend indicator. However, Figure 1 shows that the LOB can vary across features like order depths at different price levels and bid-ask spreads (BAS), even if the mid-price stays the same. Thus, relying solely on mid-price analysis may be insufficient when a deeper understanding of order book depth is necessary. This limitation is also noted in Sirignano (2019), where the authors emphasize the importance of modeling the joint distribution of bid and ask prices rather than assuming a constant bid-ask spread around the mid-price. Their approach expands mid-price classification to joint price distribution prediction, though bid and ask sizes are only used as inputs and not predicted jointly. A comprehensive view of the LOB, including bid and ask prices and their respective depths, is essential for risk management applications, such as in optimal execution, where fully capturing the LOB structure is crucial.
@@
Attention-based models like Transformer (Vaswani et al., 2017) and BERT (Devlin et al., 2019) excel at sequence-to-sequence prediction by assigning varying weights to different parts of the data, enabling them to focus on the most relevant information for more accurate predictions. These models have proven highly effective in tasks such as natural language processing (Yang et al., 2021; Galassi et al., 2020; Cui et al., 2019), speech recognition (Zeyer et al., 2018; Chorowski et al., 2015; Bahdanau et al., 2016), and time series forecasting (Shih et al., 2019; Fan et al., 2019; Du et al., 2020). Despite their success, applications of attention-based models in limit order book forecasting are still relatively limited. Recently, however, studies have highlighted their potential in this area: Arroyo et al. (2024) used survival analysis to forecast order fill probabilities, Liu et al. (2024) applied vision transformers to classify mid-price movements, and Wu et al. (2024) employed wavelet decomposition for short-term price movement prediction in the Chinese market. Nevertheless, none of these studies focused on predicting the full price and volume structure within the limit order book.
In this study, we present an attention-based approach utilizing the Spacetimeformer architecture (Grigsby et al., 2022) to address the unique challenges of limit order book (LOB) forecasting. Our method builds on previous research by introducing a customized embedding technique that effectively captures spatiotemporal features while maintaining the LOB’s structural integrity, resulting in a more detailed and accurate depiction of market dynamics. By integrating both price and volume information, this approach significantly enhances predictive accuracy compared to traditional models, showcasing the strong potential of attention-based architectures to advance forecasting in high-frequency trading contexts. These advancements highlight the critical role of attention mechanisms in identifying complex patterns within LOB data, paving the way for more sophisticated and reliable market modeling.
Our contributions are three-fold: (1) we extend the capability of attention-based models beyond mid-price prediction, enabling them to capture dynamic patterns within the entire LOB structure; (2) we implement modified spatiotemporal embeddings that effectively integrate both spatial and temporal features, offering a model generalizable to multivariate time series with compound attributes; and (3) we improve forecasting accuracy by addressing non-stationarity in high-frequency financial data using a percent-change and min-max transformation method that does not add complexity to the model.
This paper is organized as follows: Section 2 offers an overview of the dataset and experimental setup. Section 3 details the methods used to address the challenges of LOB forecasting. Section 4 presents the experimental results, comparing our model’s performance with various existing techniques. Section 5 discusses the implications of our approach, and Section 6 concludes the paper with suggestions for future research directions.
2 Data
The LOB is a real-time record of all buy and sell orders for a specific asset in financial markets. Organized by price levels, it displays the best bid (highest buy price) and best ask (lowest sell price) at any given moment, along with details like order volume and timestamps. In this section, we explore the high-frequency characteristics and spatiotemporal structure of the LOB data and outline our approach for formulating and transforming these inputs to support model development.
2.1 High-frequency dynamics and spatiotemporal correlations
Figure 2: Rolling averages of AMZN stock’s best-bid and best-ask prices across time windows ranging from 1 second to 2 hours. Higher values indicate the best-ask price being above the best-bid price. Notably, sharp changes seen around 11:00 and 12:30 are smoothed in the 1- and 2-hour windows, where shorter-term fluctuations become less prominent.
Figure 3: Bid-Ask Spread (BAS) for a selected trading day, calculated for three sample stocks: GOOG, AAPL, and AMZN.
2.1.1 High-frequency LOB
In high-frequency trading and market microstructure, traditional financial assumptions, such as stationarity and independent increments, are often considered inadequate by the real-time, high-frequency nature of the LOB. Unlike in conventional financial modeling, where these assumptions may hold over longer time horizons, the continuous flow of orders, frequent price changes, and rapid reactions by algorithmic trading result in time series data that are inherently noisy, irregular, and non-stationary. This volatility presents unique challenges in modeling and forecasting, as the traditional tools often used in lower-frequency financial data analysis do not sufficiently capture the nuanced behaviors inherent in high-frequency environments.
Figure 2 visually depicts these dynamics by showing rolling averages of AMZN stock’s best-bid and best-ask prices over various time intervals—from seconds to hours—within a single trading day. While shorter time windows capture sharper price fluctuations and greater volatility, reflecting the underlying high-frequency dynamics, longer windows smooth out these short-term variations. This visual distinction underscores how time resolution significantly impacts our understanding of data. The choice of time scale in analyzing LOB data is crucial, as high-frequency behaviors may be masked in more aggregated views, potentially overlooking the critical, rapid adjustments inherent to LOB activity.
Further complicating the analysis of LOB data is the often-used assumption of a constant bid-ask spread. This simplification is commonly applied to streamline modeling efforts; however, Figure 3 demonstrates that the bid-ask spread is far from constant in high-frequency trading. At smaller time intervals, spreads exhibit continuous and significant fluctuations, with notable changes occurring around critical times, such as market open and close. These time-dependent variations in both spread mean and volatility highlight the dynamic complexity of the LOB that a constant spread assumption cannot fully capture. While constant spread models offer mathematical tractability, they fail to represent the richness of real-time market dynamics, especially when considering the effects of microstructure in high-frequency settings (Abergel et al., 2016).
2.1.2 Spatiotemporal structure of LOB
Figure 4: Mid-price (top), level-5 bid and ask price changes (middle), and level-5 bid and ask volume changes (bottom) aligned along the same order arrivals axis. In the middle and bottom heatmaps, darker shades indicate higher absolute percentage changes, with red representing upward movements and blue representing downward movements; White areas show no change. The mid-price is shown on a raw price scale, while percentage change values are scaled for better comparability across different scales. These plots illustrate data from the first 60 order arrivals of AAPL stock at market open.
The high-frequency nature of LOB reveals complex dynamics that influence the order book movements, with multiple components contributing to these fluctuations. In a multi-level LOB, each price level and corresponding volume plays an essential role as new orders are placed, canceled, or executed, reshaping the order book’s structure and potentially impacting market behavior across different levels. These rapid and simultaneous changes reveal how the continuous interaction of market participants dynamically impacts both prices and volumes.
Figure 4 underscores this spatiotemporal relationship, presenting how mid-price movements (top panel) are shaped by changes in bid and ask prices and volumes of multiple levels. Price shifts in the multi-level structure reflect the cumulative effect of variations in both price and volume, which tend to be interconnected. In the middle and bottom heatmaps, we observe that large portions of the LOB remain static, showing 0% change, meaning that not all levels are altered with each order event. This sparse activity, where only specific levels undergo changes, suggests that most of the LOB remains stable between updates, with shifts happening intermittently across some levels. However, when a substantial movement occurs in a certain direction, it often influences neighboring levels, driving future movements in a similar direction. Notably, these movements often occur concurrently across various attributes, including price and volume, suggesting a compounded effect across the LOB. This sparse yet correlated activity provides valuable insight into how market depth and order flow dynamically shape the LOB across multiple levels.
These observations emphasize the need for a modeling approach that can accommodate the high-frequency, time-varying, and spatiotemporal properties of LOB data. In this light, a data-driven method, rather than relying on traditional theoretical assumptions, is better suited to capture dependencies that evolve over short intervals. Adopting such an approach allows for a more accurate representation of the dynamics in the LOB, providing insights into market behavior that can inform effective decision-making in high-frequency trading contexts.
@@ 2.2 Data structure and preparation
We utilize the LOBSTER dataset (Huang & Polak, 2011), which offers real-world LOB data, including both message and order book information for selected stocks. We specifically focus on five major technology companies: AAPL, GOOG, INTC, MSFT, and AMZN. The sample covers the trading day of June 21, 2012, from the market opening at 9:30 am to the closing at 4:00 pm, capturing between 300,000 to 600,000 snapshots per stock, all recorded with millisecond time resolution at Level-5 of the order book. The order book file tracks the evolution of the LOB, while the message file logs the events that trigger updates, with timestamps ranging from milliseconds to nanoseconds.
@@
2.3 Input transformation and scaling
@@
The performance of various transformation methods is compared in a later section, specifically in Table 1 within Section 4.
@@ 3.1 Compound multivariate embedding
$\text{Attention}(Q,K,V)\in\mathbb{R}^{L_{x}\times d}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V,$ (3)
@@ 4 Results
Table 1: Performance metrics calculated for each input transformation calculated for each input transformation across the top-5 price levels. Mean Squared Error (MSE) and Mean Absolute Error (MAE) are provided for each method. All values have been descaled and are presented in their original units of dollars.
Input transformation Percent-change Min-max Percent-change + Minmax
MSE 4.5447 0.0660 0.0464
MAE 1.3225 0.1469 0.1260
Table 2: Prediction results for sample test data: Mean Squared Error (MSE) and Mean Absolute Error (MAE) are computed for each method across selected dimensions. The mid-price is determined by averaging the best bid and ask prices. Loss metrics are provided for mid-price, price, and volume on level-5 limit order book (LOB) data. Forecasting loss refers to MSE losses from all five levels of prices and volumes, while structure loss quantifies the violations of the ordinal structure. Total loss represents the weighted sum of the forecasting loss and structure loss. All values are calculated after scaling.
@@
mid-price MSE 0.0125 0.0124 0.0124 0.0122 0.0120
MAE 0.1501 0.1466 0.1520 0.1425 0.1388
price* MSE 0.0026 0.0026 0.0027 0.0025 0.0025
MAE 0.0310 0.0303 0.0332 0.0293 0.0288
volume** MSE 0.0105 0.0118 0.0109 0.0105 0.0106
MAE 0.0510 0.0541 0.0533 0.0515 0.0504
Forecasting loss 0.0065 0.0072 0.0068 0.0065 0.0066
Structure loss 0.2430 0.2624 0.8836 0.5774 0.1480
Total loss 0.0090 0.0098 0.0157 0.0123 0.0080
@@
To assess the impact of input transformations on prediction accuracy, we first examine the results in Table 1. Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics are reported for different transformations applied to multi-level price data. The combined approach, using both percent-change and min-max scaling, achieves the lowest error rates with an MSE of 0.0464 and an MAE of 0.1260, outperforming each transformation used individually. These values are presented in the raw scale to allow a consistent comparison of transformation effects.
Percent-change normalization, a standard transformation in financial time series, enforces stationarity in input sequences—an essential condition in time series forecasting. In machine learning, non-stationarity of time series can lead to unreliable predictions, as trends or shifts may skew the model’s focus away from the underlying fluctuations. By transforming prices into percent changes, the model can emphasize relative movements around a baseline, capturing meaningful patterns without the distraction of directional trends.
Min-max scaling, on the other hand, is well-suited for multivariate data where different scales are present, as in order volume and price levels. Volume data often has higher variance than price data, and min-max scaling helps optimize learning across price levels by constraining these variations. This approach prevents extreme volume values from overwhelming the model’s attention, enabling it to respond effectively to both price and volume information.
Our experimental results demonstrate that combining percent-change normalization and min-max scaling maximizes the benefits of each transformation. By stabilizing the price series for stationarity and managing volume variations within a fixed range, this combined approach achieves more accurate and consistent predictions across multi-level price data, resulting in enhanced model performance and reliability.
@@
We now present the prediction results for test data, with a focus on various performance metrics discussed in Section 3.2. Our experiments compare various sequence-to-sequence prediction models of prior work, including a linear autoregressive model (Linear), a standard LSTM network (Schmidhuber et al., 1997) (LSTM), a transformer with temporal attention (Temporal) as introduced in Informer (Zhou et al., 2021), and a spacetimeformer model (Spacetime) that incorporates both spatial and temporal attention (Grigsby et al., 2022). We also evaluate the performance of our attention-based model with compound multivariate embedding (Compound) in comparison to the other techniques.
Table 2 compares overall prediction results of various models on level-5 LOB data, with the compound attention-based model (ours) showing the best overall performance. This model achieved the lowest error in all price-related metrics, including mid-price and multi-level prices, outperforming both traditional time series models and other attention-based models. Notably, the compound model excelled in maintaining ordinal structure with the lowest structure loss, leading to the lowest total loss, indicating a strong balance of accuracy and structural consistency across mid-price, price, and volume forecasting. This highlights the method’s superior ability to capture structural dependencies through the compound multivariate embedding process. Although metrics calculated from mid-price, multi-level prices and volumes are recorded for reference, they are not used as loss functions.
@@
Figure 9 and Figure 10 provide visual comparisons of the actual versus predicted price and volume movements for the level-5 LOB data of AMZN. We selected the first window of the test period for comparison. In Figure 9, the price forecasting results demonstrate that the compound multivariate embedding method (B) closely tracks the downward trend of actual price movements (A), offering more accurate predictions than the temporal (C) and spatiotemporal (D) embedding methods. While both our approach and the spatiotemporal embedding effectively capture the overall ordinal structure of multi-level prices, the temporal embedding fails to preserve the ordinal relationships between dimensions. This aligns with the concepts of the embedding methods, as the temporal embedding processes multiple time series without incorporating variable embeddings, resulting in structural violations.
Similarly, in Figure 10, the volume forecasting results indicate that the compound multivariate embedding method (B) produces a snapshot of volume distributions that align closely with the actual order sizes (A). Differences among the attention-based models are less pronounced for volume predictions than for price forecasts. Additional examples of outputs from all models, including both time series and attention-based approaches, are provided for other stocks, with results shown in Figures S1, S2, S3, S4, S5, and S6,.
@@
5 Discussions
Our approach to forecasting the entire LOB is fully data-driven, allowing for predictions of both multi-level prices and their respective order sizes at each level. This multi-level perspective expands our analytical horizon beyond mid-price forecasting, where practical applications are often limited. By capturing the granular structure of price levels, our model better reflects real market dynamics, providing a more actionable foundation for order execution.
In addition to price information, market depth—represented by the order sizes across levels—plays a significant role in influencing the LOB’s rapid fluctuations. Events such as the addition, cancellation, or execution of large orders lead to priority shifts and reordering within levels, causing the LOB to dynamically adjust as orders enter or exit different price tiers. Our comprehensive, data-driven approach enables a more accurate depiction of market behavior by integrating the full scope of LOB information. Due to computational resource limitations, our experiments were conducted on the top five levels of the LOB. However, our method is extendable, allowing for deeper depth levels and broader time horizons.
While our model effectively captures dependencies by focusing on multi-level price and volume data alone, one limitation lies in its inability to identify the precise causes of these fluctuations—whether due to new orders, cancellations, executions, or modifications. Using an attention-based architecture, the model adapts in real time to shifts in market context, accurately tracking evolving dependencies within the LOB. However, future work could enhance this by incorporating the specific types of order events, allowing the model to differentiate between sources of LOB changes and better interpret their unique effects on LOB dynamics.
In practice, the ability to forecast multi-level LOB changes can significantly enhance strategies for optimal execution of large orders by simulating the LOB’s response based on current data. This approach allows traders to anticipate how their trades might affect market depth, spread, and price dynamics, enabling them to adjust order placement to reduce slippage and minimize market impact. By leveraging multi-level insights, they can optimize trade size and timing to prevent adverse price moves and execute in a way that preserves liquidity.
@@ 6 Conclusions
In this study, we addressed the challenge of forecasting multi-level LOB data, a task that often exceeds the capabilities of conventional time-series forecasting models. Our primary contribution is the development and implementation of a compound multivariate embedding method within advanced sequence-to-sequence models. This method effectively captures the complex interdependencies among various LOB attributes, including order types, features, and levels. The empirical results demonstrate that our approach consistently outperforms existing multivariate forecasting methods, achieving the lowest forecasting errors while maintaining the ordinal structure of the LOB. These findings highlight the effectiveness of our method for financial time-series forecasting, particularly in the context of high-frequency trading. As a direction for future work, our method could be extended to address optimal execution problems, with further optimization for real-time forecasting scenarios to enhance its practical applicability in dynamic trading environments.
S1 Supplementary Material
Figure S1: Price forecasting of a limit order book for GOOG stock: The graphs predict level-5 prices 2 minutes ahead, based on the preceding 10 minutes of context, with the prediction period starting on the right side of the dotted line. Panel A (top left) shows the actual price movements, while panels B, C, and D display the predicted price movements using five forecasting methods: compound multivariate embedding (B), temporal embedding (C), spatiotemporal embedding (D), linear model (E), and LSTM network (F). Multi-level prices, excluding the best prices, are represented in sequential grey lines.
Figure S2: Price forecasting of a limit order book for MSFT stock: The graphs predict level-5 prices 2 minutes ahead, based on the preceding 10 minutes of context, with the prediction period starting on the right side of the dotted line. Panel A (top left) shows the actual price movements, while panels B, C, and D display the predicted price movements using five forecasting methods: compound multivariate embedding (B), temporal embedding (C), spatiotemporal embedding (D), linear model (E), and LSTM network (F). Multi-level prices, excluding the best prices, are represented in sequential grey lines.
Figure S3: Price forecasting of a limit order book for AAPL stock: The graphs predict level-5 prices 2 minutes ahead, based on the preceding 10 minutes of context, with the prediction period starting on the right side of the dotted line. Panel A (top left) shows the actual price movements, while panels B, C, and D display the predicted price movements using five forecasting methods: compound multivariate embedding (B), temporal embedding (C), spatiotemporal embedding (D), linear model (E), and LSTM network (F). Multi-level prices, excluding the best prices, are represented in sequential grey lines.
Figure S4: Volume forecasting of a limit order book snapshot for GOOG stock: The graphs forecast level-5 volumes for the first snapshot of the testing period based on the preceding 10 minutes of context. Panel A shows the actual order sizes, while panels B, C, and D display the predicted volume snapshots using five forecasting methods: compound multivariate embedding (B), temporal embedding (C), spatiotemporal embedding (D), linear model (E), and LSTM network (F).
Figure S5: Volume forecasting of a limit order book snapshot for MSFT stock: The graphs forecast level-5 volumes for the first snapshot of the testing period based on the preceding 10 minutes of context. Panel A shows the actual order sizes, while panels B, C, and D display the predicted volume snapshots using five forecasting methods: compound multivariate embedding (B), temporal embedding (C), spatiotemporal embedding (D), linear model (E), and LSTM network (F).
Figure S6: Volume forecasting of a limit order book snapshot for AAPL stock: The graphs forecast level-5 volumes for the first snapshot of the testing period based on the preceding 10 minutes of context. Panel A shows the actual order sizes, while panels B, C, and D display the predicted volume snapshots using five forecasting methods: compound multivariate embedding (B), temporal embedding (C), spatiotemporal embedding (D), linear model (E), and LSTM network (F).